OpenStudy (lukecrayonz):

log(16x+2)+log(20x-2)=log(319x^2+9x-2) log(16x+2)+log(20x-2)=log(319x^2+9x-2) @Mathematics

OpenStudy (lukecrayonz):

This is supposed to be a puzzle for extra credit, but I already solved the puzzle, I just need to show work. So the answer is x=2.

OpenStudy (anonymous):

solve \[(16x+2)(20x+2)=319x^2+9x-2\] for x

OpenStudy (anonymous):

good work satellite.........hey can i ask one thing from u @ satellite

OpenStudy (lukecrayonz):

That didnt work satellite :S

OpenStudy (anonymous):

typo there, should be \[(16x+2)(20x-2)=319x^2+9x-2\] http://www.wolframalpha.com/input/?i=%2816x%2B2%29%2820x-2%29%3D319x%5E2%2B9x-2

OpenStudy (anonymous):

you get two values for x, -1 and 2, but your only answer is 2 because if x = -1 then 20x -2 = -22 and you cannot take the log of a negative number

OpenStudy (anonymous):

x^2 - x -2 = 0 solve for x

OpenStudy (anonymous):

don't forget to drop the "-1" solution @sheg, ask away